OpenStudy (anonymous):

For the circuit below, each branch has equal resistance. What value of resistance (R) will produce an equivalent resistance of 33.3 Ω?

OpenStudy (anonymous):

\[V, R!,R2, are the resistors on the circuit and they are blank\]

OpenStudy (anonymous):

V,R1,R2 are the resistors on the circuit and the are blank

OpenStudy (radar):

Besides telling you they are equal, do they have a value assigned to any of them? Or is the voltage and current given? You are missing information needed to solve this problem.

OpenStudy (anonymous):

are the resistors in series or parallel

OpenStudy (anonymous):

sorry they are parallel and they have no value I couldn't upload a picture but are two circuits R1, R2 and Voltage and there is no current given

OpenStudy (anonymous):

each resistor is 100 ohms

OpenStudy (anonymous):

(1 )/( R_T )=1/R_1 +1/R_2 +1/R_3

OpenStudy (anonymous):

.is it three resistor or two. i did for three.

OpenStudy (anonymous):

its only two

OpenStudy (anonymous):

i am not getting to see the circuit if it two then each is around 67 ohms.

OpenStudy (anonymous):

is there a way i can upload it here

OpenStudy (anonymous):

do you understand the formular i put there. it is a given if two resistor in parallel with the same value you divide by two if it 3 with the same value divide by 3.

OpenStudy (anonymous):

what subject is that?

OpenStudy (anonymous):

yes i understand the forumla but my biggest issue was finding the value of the two resistors because all was given was the equivalnent resistance of 33.3 ohms this is for my ECET class at devry.

OpenStudy (anonymous):

but its physics it just sucks because my professor is really not that good at responding back and the book that we have is really not great for sshowing you how to do things

OpenStudy (anonymous):

ok.

OpenStudy (radar):

We are making some assumptions. For example if the two resistors were 400 ohms the added resistor to make an equivalent amount of resistance of 33.3 would be a 40 Ohms. Check it out

OpenStudy (anonymous):

yes rada that it we want to add a 3rd resistor but we only given two in parallel and the total. we were not ask what value should we add to get 33.3ohms.

OpenStudy (anonymous):

i tried to jput it as my profile picture lol but its to small

OpenStudy (radar):

The assumption was that the two were 100 Ohm and a 3rd 100 ohm would indeed be a 100 ohm, but there are many other possibilities.

OpenStudy (radar):

I thought we were given a circuit with two branches with resistors and requested what would be the value of a resistor to connect to a third branch to make the equivalent resistance 33.3 Ohms.

OpenStudy (anonymous):

_________________ | | | V| R1{ R2 { | | | | | | | | | |______ |_________ |

OpenStudy (radar):

I guess if the third resistor had to equal the other two, then 100 can be the only answer

OpenStudy (anonymous):

ok its not perfect but thats how it looks

OpenStudy (radar):

Does it say that the three resistors have to be equal?

OpenStudy (anonymous):

no it says that the two resistors have to be equal

OpenStudy (anonymous):

I have one that says that is the same as this one that has three branchs that have to be equal but for this problem it only request two branches to be equal

OpenStudy (anonymous):

this is where i kept getting confused because the book doesn't explain this and I couldn't figure out what to do.

OpenStudy (radar):

Well 100 for each will work, so will 400 for the two and 40 for the third.

OpenStudy (anonymous):

if you have any more question email me at lavicrep5@hotmail.com

OpenStudy (anonymous):

radar read the question again. i think we might confuse the person.

OpenStudy (anonymous):

see thats what i thought last night about 100 for both but then when you add them up and then divide them they don't come out to 33.3 if i divide 3 from 100 I come up with 33.3

OpenStudy (radar):

I've read it agaIn and I find it confusing? Is the final circuit to be 2 or 3 resistors?

OpenStudy (radar):

The resistors are connected in the parallel configuration (branches rather then series.

OpenStudy (radar):

The total resistance of parallel connected resistors will be less than the smallest resistor.

OpenStudy (anonymous):

just 2

OpenStudy (radar):

Also the total resistance is the recripocal of the sum of the reciprocal values.

OpenStudy (anonymous):

yes so i thought that if i divided 33.3 that would give me my answer for the two resistors but i was not sure about that

OpenStudy (radar):

O.K then they would be 66 2/3 ohms each. Like resistors when connected parallel, the total can be computed by dividing the resistor value by the number of resistors.

OpenStudy (anonymous):

ok. i think that the final result

OpenStudy (radar):

prcival, did you find the wording of the problems sort of confusing?

OpenStudy (anonymous):

yep it is confusing. you get to know your circuit to sort it out. it was left for too many interpretations.

OpenStudy (anonymous):

tell me about i had to reread it a couple of times to make sure i was looking at it right

OpenStudy (anonymous):

thank you for helping me because i was completly lost

OpenStudy (anonymous):

so i would just do the same thing for the three resistors correct

OpenStudy (anonymous):

percival I had sent you a pic of the circuit did you receive it

OpenStudy (anonymous):

yes. do you have more?

OpenStudy (anonymous):

yes i have another one just like the one we did earlier but its with three resistors but its the same problem

OpenStudy (anonymous):

so i made all of them 100 then divde by three is that correct

OpenStudy (anonymous):

ok then in that case you multiply the total by 3. remind you if they are is series you cannot multiply youhave to divide and that only work when the rsistor values are the same.

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

cool i so understand now you have been a life saver

OpenStudy (anonymous):

you welcome feel free to email me questions.

OpenStudy (anonymous):

i will you guys just don;t know how relieved I am now to finally understand this

OpenStudy (anonymous):

Answer is 66.6 ohms. If you have 2 resistors with the same value,for a sample;R1=50 and R2=50 ,just divide one on half,50/2=25. In your problem 66.6/2=33.3 ohms. Or you can do this way R1xR2/R1+R2=66.6x66.6/66.6+66.6=33.3 Ohms.

OpenStudy (anonymous):

Answer is 66.6 ohms. If you have 2 resistors with the same value,for a sample;R1=50 Ohm and R2=50 Ohm ,just divide one on half,50/2=25 Ohm ,total resistance. In your problem 66.6/2=33.3 ohms.R1=66.6 Ohm,R2=66.6 Ohm Or you can do this way R1xR2/R1+R2=66.6x66.6/66.6+66.6=33.3 Ohms is the total resistance in the circuit.